OpenStudy (anonymous):

You are having a meeting with the CEO of a soda company. You have interpreted the number of cans of soda produced versus profit as the function P(x) = x4 + 2x3 + 6x2 - 3x - 7. Describe to the CEO what the graph looks like. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0). @satellite73 I need your help plz :)

OpenStudy (anonymous):

really? here is the graph http://www.wolframalpha.com/input/?i=+x4+%2B+2x3+%2B+6x2+-+3x+-+7

OpenStudy (anonymous):

you sure you didn't make a typo there right? the zeros are impossible to find without technology here

OpenStudy (anonymous):

for the "end behavior" you see that as \(x\to \infty\) you have \(p(x)\to \infty\)

OpenStudy (anonymous):

in other words, the more cans of soda you make the more money you make

OpenStudy (anonymous):

i'm sure it isn't a typo i copied it word for word

OpenStudy (anonymous):

since you cannot make negative cans of soda, we should restrict the domain of this function to \(x>0\)

OpenStudy (anonymous):

it looks ilke the graph crosses the x axis at just less than \(x=1\) so if you make only one can of soda you still make a profit

OpenStudy (anonymous):

okay I'll send you my answer and please tell me if its right or not

OpenStudy (anonymous):

can u fan me so i cand send u an inbox of my answer?

OpenStudy (anonymous):

this is an amazingly idiotic question if you make only 10 cans of soda your profit is \[p(10)=\$12,563\]

OpenStudy (anonymous):

i really need help im stuckk :/ I dont know what to doo!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets go slow and do what it asks you to do

OpenStudy (anonymous):

\[P(x) = x^4 + 2x^3 + 6x^2 - 3x - 7 \] is supposed to be your profit on \(x\) cans of soda

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Describe to the CEO what the graph looks like for values of \(x\) that are greater than zero, the graph grows very quickly

OpenStudy (anonymous):

i will send you a picture it almost looks like a parabola, but not quite http://www.wolframalpha.com/input/?i=+x4+%2B+2x3+%2B+6x2+-+3x+-+7+domain+0..10

OpenStudy (anonymous):

you can see that if you make only ten cans of soda you make over \(\$12,000\) i didn't write this question, i call them as i see them

OpenStudy (anonymous):

the end behaviour: the more cans of soda you make, the bigger your profit

OpenStudy (anonymous):

where the company will break even (where P(x) = 0) for this one \(p(x)=0\) for \(x\) just about 1, which means if you make one can of soda you will make a profit

OpenStudy (anonymous):

no point in asking again if you copied this correctly, i believe you but it is rather silly i wish i could make this kind of money

OpenStudy (anonymous):

there is not really anything else to say i did fan you if you need to send me a message

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

is that answer complete enough for you? not sure what else to say

OpenStudy (anonymous):

its kind of confusing for me its really hard to understand they always give us these really dumb question that we've never even went through during class and just expect you to know whatever they give u out of the blues.

OpenStudy (anonymous):

ok your answer makes sense in terms of the graph, but no in terms of the question the question has x as the number of cans of soda made you can't really make negative cans of soda

OpenStudy (anonymous):

but you are doing the best you can for the tools that you got, so go ahead with your answer you also are supposed to say what the "end behaviour" is, so include the fact that the more cans of soda you make, the more money you make

OpenStudy (anonymous):

can u help me out by adding that cuz i have no idea what else to say if you don't want to i understand :)

OpenStudy (anonymous):

just say what i just wrote as \(x\to \infty, p(x)\to \infty\) so the more cans of soda you make, the higher your profit you can copy that verbatim

OpenStudy (anonymous):

that should be good enough

OpenStudy (anonymous):

thank you so much i appreciate it! :)

OpenStudy (anonymous):

i have one more i'm stuck on

OpenStudy (anonymous):

lets get to it then

OpenStudy (anonymous):

thankss :)

OpenStudy (anonymous):

Your friend hands you a graph of the engine power statistics of a race car. He says, “I know this graph is f(x) = –3(x + 4)6 – 8 but I can’t remember how it is related to the graph of x6.” Explain to your friend how the graph f(x) is a translation of the graph x6.

OpenStudy (anonymous):

dreaded question marks again let me refresh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x)=-3(x+4)^6-8\]

OpenStudy (anonymous):

how does this translate from \[y=x^6\]

OpenStudy (anonymous):

1) from \(y=x^6\) to \(y=(x+4)^6\) the graph shifts 4 units to the LEFT (write these down)

OpenStudy (anonymous):

okay :) @satellite73

OpenStudy (anonymous):

2) from \(y=(x+4)^6\) to \(y=-3(x+4)^6\) the graph rotates (flips) about the \(x\) axis and becomes narrower by a factor of 3

OpenStudy (anonymous):

3) from \(y=-3(x+4)^6\) to \(y=-3(x+4)^6-8\) the graph shifts DOWN 8 units

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

here is the graph on the interval \((-4,2)\) http://www.wolframalpha.com/input/?i=-3%28x%2B4%29^x-8+domain+-4..2

OpenStudy (anonymous):

alright @satellite73

OpenStudy (anonymous):

hope you are done